40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Entry  Mon Oct 9 01:29:26 2017, gautam, Update, LSC, DRMI Nosie Budget v3.0 7x
    Reply  Mon Oct 9 11:55:01 2017, Koji, Update, LSC, DRMI Nosie Budget v3.0 
       Reply  Mon Oct 9 22:18:34 2017, gautam, Update, LSC, AS55Q Dark Noise AS55Q_dark.pdf
          Reply  Tue Oct 10 22:04:06 2017, rana, Update, LSC, AS55Q Dark Noise 
          Reply  Wed Oct 11 14:42:03 2017, gautam, Update, LSC, AS55Q Dark Noise AS55Q_darkNoises.pdf
             Reply  Wed Oct 11 19:31:32 2017, gautam, Update, LSC, AS55Q Dark Noise AS55Q_darkNoises.pdf
                Reply  Thu Oct 12 01:50:11 2017, gautam, Update, LSC, AS55Q Dark Noise AS55_dark.png
                   Reply  Thu Oct 12 12:17:28 2017, gautam, Update, LSC, AS55Q Dark Noise D040179_AS55_40m.pdf
                      Reply  Fri Oct 13 12:26:12 2017, gautam, Update, LSC, AS55Q Dark Noise AS55_DemodNoises.pdf
                         Reply  Mon Oct 16 16:01:04 2017, gautam, Update, LSC, AS55Q Dark Noise DemodBoardwOscope.pdfDemodBoardwOscope_ASD.pdf
                            Reply  Tue Oct 17 19:31:53 2017, gautam, Update, LSC, AS55Q Dark Noise D040179_TFs.pdfAS55_DemodNoises.pdf
    Reply  Tue Nov 7 17:45:05 2017, gautam, Update, LSC, DRMI Nosie Budget v3.1 SRCL_MICH_whitenoise_tf.pdfPRCL_MICH_whitenoise_tf.pdfMICH_aux.pdfC1NB_disp_40m_MICH_NB_2017-10-08.pdf
       Reply  Tue Nov 7 22:56:21 2017, gautam, Update, LSC, DRMI locking recovered DRMI_7Nov20178.png
          Reply  Wed Nov 8 00:28:16 2017, gautam, Update, LSC, Laser intensity coupling measurement attempt PSL_RIN.pdf
       Reply  Wed Nov 8 09:37:45 2017, rana, Update, LSC, DRMI Nosie Budget v3.1 
          Reply  Wed Nov 8 09:59:12 2017, gautam, Update, LSC, DRMI Nosie Budget v3.1 
             Reply  Wed Nov 15 01:37:07 2017, gautam, Update, LSC, DRMI low freq. nosie improved C1NB_disp_40m_MICH_NB_2017-11-15.pdfOL_A2L_couplings.pdf
                Reply  Thu Nov 16 00:53:26 2017, gautam, Update, LSC, DRMI noise sub-budgets C1NB_a2l_40m_MICH_NB_2017-11-15.pdfC1NB_aux_40m_MICH_NB_2017-11-15.pdf
Message ID: 13428     Entry time: Wed Nov 15 01:37:07 2017     In reply to: 13416     Reply to this: 13431
Author: gautam 
Type: Update 
Category: LSC 
Subject: DRMI low freq. nosie improved 

Pianosa just crashed and ate my elog, along with all the DTT/Foton windows I had open, so more details tomorrow... This workstation has been crashing ~once a month for the last 6 months.

Summary:

Below ~100Hz, the hypothesis is that the BS oplev A2L contribution dominates the MICH displacement noise. I wanted to see if I could mitigate this my modifying the BS Oplev loop shape.

Details:

  • Swept sine TF measurements suggested that the BS A2L contribution is between 10-100x that of the ITM A2L
  • The Oplev loop shape for BS is different from ITMs - specifically, there is a Res-gain centered at ~3.3 Hz. The low frequency ~0.6Hz boost filter present in the ITM Oplev loops was disengaged for the BS Oplves.
  • I turned off the BS OL loops and looked at error signal spectra - didn't seem that different from ITM OL error signals, so I decided to try turning off the res-gain and engage the 0.6Hz boost.
  • This change also gave me much more phase at ~6Hz, which is roughly the UGF of the loop. So I put in another roll-off low pass filter with corner frequency 25Hz. 
  • This worked okay - RMS went down by ~5x (which is even better than the original config), and although the performance between ~3 and 10Hz is slightly worse than with the old combination,this region isn't the dominant contribution to the RMS. PM at the upper UGF is ~30degrees in the new configuration. 
  • I wanted to give DRMI locking a shot with the new OL loop - expectations were that the noise between 30-100Hz would improve, and perhaps the engaging of de-whitening filters on BS would also be easier given the more severe roll-off at high-frequencies.
  • Attachment #1 shows the NB for tonights lock. All MICH optics had their coil drivers de-whitened, and all the LSC PDs were whitened for this measurement.
  • I've edited the NB code to make the A2L calculation more straightforward, I now just make the coupling 1/f^2 and give the function a measured overall gain, so that this curve can now be easily added to all future NBs. I've also transcribed the matlab funciton used for parsing Foton files into python, this allows me to convert the DQ-ed OL error signals to control signals. Will update git with changes.

Remarks:

  1. MICH noise has improved by ~2x between 40-80Hz.
  2. Not sure what to make of the broad hump around 60Hz - scatter shelf?
  3. There is still unexplained noise below 100Hz, the A2L estimate is considerably lower than the measured noise.
  4. We are still more than an order of magnitude away from the estimated seismic noise floor at low frequencies (but getting closer!).

I've been banging my head against optimal loop shaping, with the OL loop as a test-case, without much success - as was the case with coating PSO, the magic is in smartly defining the cost function, but right now, my optimizer seems to be pushing most of the roots I'm making available for it to place to high frequencies. I've got a term in there that is supposed to guard against this, need to tweak further...


Attachment #2: Eye-fits of measured OL A2L coupling TFs to a 1/f^2 shape, with the gain being the parameter "fitted". I used these value, and the DQ-ed OL error signal in lock, to estimate the red curve labelled "A2L" in Attachment #1. The dots are the measurement, and the lines are the 1/f^2 estimates.

Attachment 1: C1NB_disp_40m_MICH_NB_2017-11-15.pdf  465 kB  | Hide | Hide all
C1NB_disp_40m_MICH_NB_2017-11-15.pdf
Attachment 2: OL_A2L_couplings.pdf  93 kB  Uploaded Wed Nov 15 12:00:52 2017  | Hide | Hide all
OL_A2L_couplings.pdf
ELOG V3.1.3-